Following information is extracted from the books of XYZ Traders on 31st December
2018.
Particulars Rs. Particulars Rs.
Salesman salary 60,500 Sales revenue 121,000
Plant & Machinery 750,000 Land & Building 840,000
Furniture and Fixtures 170,000 Long term loans 45,000
Goodwill 100,000 Patent rights 150,000
Gross profit 24,000 Commission received 20,000
Copy rights 220,000 Salaries of marketing staff 172,000
Sundry creditors 25,000 Mark up on loan 14,000
Salaries of office staff 160,000 Advertising expenses 63,000
Electricity charges of office
building
95,000 Short term loans 33,000
Rent of office building 60,000 Taxes paid 31,000
Required:
Based on the above information of XYZ Traders, you are required to calculate the
amount of:
1. Tangible Assets
2. Intangible Assets
3. Administration Expenses
4. Selling Expenses
1
Expert's answer
2019-05-27T11:49:58-0400

1. Tangible Assets

Tangible assets refers to the physical and measurable assets in company which are used in operations. Examples include property, plant, and equipment, land, machinery, buildings, and inventory. For XYZ Traders, the tangible assets listed are Plant & Machinery (750,000), Land & Building (840,000), and Furniture and Fixtures (170,000).

Thus,

Tangible assets= 750,000 + 840, 000 + 170,000

                      =1,760,000

2. Intangible Assets

Intangible assets are those assets of a company which are non-physical. Examples are copyrights, patents, goodwill, trademarks, and brand recognition. For XYZ traders, intangible assets are goodwill (100,000), patent rights (150,000), and copy rights (220,000).

Thus,

Intangible assets= 100,000 + 150,000 + 220,000

                          =470,000

3. Administration Expenses

Administration expenses refers to those expenses which an organization incurs and is not directly tied to a particular function such as sales, production, or manufacturing. Examples are interest expense, debt placement costs, bank fees, auditor fees, senior executives’ salaries, and cost of general services like accounting. For XYZ traders’ administrative expenses are salaries of office staff (160,000), rent of office building (60,000), and electricity charges of office building (95,000).

Thus,

Administrative expenses= 160,000+ 60,000 + 95,000

                                    =315,000

4. Selling Expenses

Selling expenses refers to the costs which are associated with marketing, distributing, and selling a product or service. Examples are shipping costs, logistic costs, insurance costs, advertising costs, social media spending, website maintenance costs, wages, out-of-pocket expenses, and commissions. For XYZ traders, the selling expenses are advertising expenses (63,000), Commission received (20,000), salaries of marketing staff (172,000), and salesman salary (60,500).

Thus,

Selling expenses= 63,000 + 20,000 + 172,000 + 60,500

                        =315,500